OpenStudy (anonymous):

solve for x x=-3|x+7|=-12 A. x=5/3, x=-19/3 B. x=-3, x=-11 C. x=-3, x=11 D. no solution

OpenStudy (johnweldon1993):

We divide both sides by -3 \[\large |x + 7| = 4\] Now we break this up into 2 equations \[\large x + 7 = 4\] and \[\large x + 7 = -4\]